if the complement of angle D is 27 degrees, then what is the measure of the supplement of angle D
Answer by drj(1380)

If the complement of angle D is 27 degrees, then what is the measure of the supplement of angle D

Step 1. Complementary angles means two angles add up to 90 degrees.

Step 2. Let D be the angle and 90-D be its complement.

Step 3. 27+D=90 or D=90-27=63

Step 4. Supplementary angles means two angles add up to 180 degrees.

Step 5. Let 180-D be the supplementary angle of D or 180-63=117

Step 6. ANSWER: The supplementary angle of D is 117 degrees.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J